OpenStudy (chris215):

If f is a function such that lim (f(x)-f(a))/(x-a)=7 as x->a, then which of the following statements must be true?

OpenStudy (chris215):

f(a) = 7 f(x) is continuous at x = 7 the linear approximation for f(x) at x = a is 7 f '(a) = 7

zepdrix (zepdrix):

Here is one of the limit definitions for the derivative.\[\large\rm f'(a)=\lim_{x\to a}\frac{f(x)-f(a)}{x-a}\]

zepdrix (zepdrix):

So if we have this:\[\large\rm f'(a)=\color{orangered}{\lim_{x\to a}\frac{f(x)-f(a)}{x-a}=7}\]What does that tell you?

OpenStudy (chris215):

no we arent

zepdrix (zepdrix):

Strange.. because it's clearly the last option. But the second also seems to be true. Differentiability implies continuity. So our function should be continuous at x=7 since the derivative exists there.

OpenStudy (chris215):

hm ok

OpenStudy (zarkon):

"Differentiability implies continuity. So our function should be continuous at x=7 since the derivative exists there." "f(x) is continuous at x = 7" Might not be true. What we do know is that \(f(x)\) is continuous at \(x=a\) therefore the only answer must be \(f '(a) = 7\)

zepdrix (zepdrix):

Oh continuous at x=a, yes yes yes c: Another derp moment. Zark to save the day.